Solve
(4x + 3)^2 = 5

Answer by richwmiller(17219)

It is a parabola .
4x+3=+\- sqrt(5)
4x=-3+\- sqrt(5)
x=-3/4+\- sqrt(5)/4 or
x=1/4(-3+\- sqrt(5))